# Iterative Compositional Data Generation for Robot Control

### v1

@@ 5 Related Work
Compositional Generalization in Robotics and RL   In robotics, compositional generalization has been pursued through a variety of mechanisms. Some approaches introduce modularity or architectural biases aimed at composing semantic units such as instructions or high-level skills (kuo2020deepcompositional; wang2023programmatically; xu2018ntp; devin2019planarithmetic). Other work targets the control layer directly, designing modular, factorized, or entity-centric policy architectures that encourage reuse of behavioral components across tasks (mendez2022modular; zhou2025policy; devin2017learning). Some approaches seek to automatically identify and decompose policies into functional modules (yang2020multitask; goyal2021recurrent; mittal2020learning). A complementary direction exploits scene-centric formulations that use structured object-relational representations to compose low-level visuomotor skills in novel physical configurations (qi2025composebyfocus).
@@
Generated Data in Robotics and RL    Synthetic data has become a key idea for scaling robotic learning. One line of work expands imitation datasets through trajectory-level augmentations, where expert demonstrations are perturbed, resampled, or regenerated to increase coverage (Mandlekar2023; Jiang2025DexMimicGen; RoCoDA2024; RoboGen2023). Although these methods enrich demonstration sets, they remain constrained to variations of the same underlying tasks without expanding into new compositions.
@@
Another orthogonal direction emphasizes visual augmentation, including render-driven and vision-only pipelines that procedurally generate synthetic video datasets (Singh2024; Bonetto2023; Yu2024VisualParkour; Re3Sim2024; Real2Render2024), as well as generative and diffusion-based methods that augment images while holding actions constant (chen2023genaug; yu2023scaling). While effective for increasing visual diversity, these approaches do not provide transition-level data reflecting novel task semantics.
@@ Acknowledgments
DSB acknowledges support from the Center for Curiosity.

### v2

@@ 5 Related Work
Compositional Generalization in Robotics and RL   In robotics, compositional generalization has been pursued through a variety of mechanisms. Some approaches introduce modularity or architectural biases aimed at composing semantic units such as instructions or high-level skills (xu2018ntp; devin2019planarithmetic; kuo2020deepcompositional; wang2023programmatically). Other work targets the control layer directly, designing modular, factorized, or entity-centric policy architectures that encourage reuse of behavioral components across tasks (devin2017learning; mendez2022modular; zhou2025policy). Some approaches seek to automatically identify and decompose policies into functional modules (yang2020multitask; mittal2020learning; goyal2021recurrent). A complementary direction exploits scene-centric formulations that use structured object-relational representations to compose low-level visuomotor skills in novel physical configurations (qi2025composebyfocus).
@@
Generated Data in Robotics and RL    Synthetic data has become a key idea for scaling robotic learning. One line of work expands imitation datasets through trajectory-level augmentations, where expert demonstrations are perturbed, resampled, or regenerated to increase coverage (Mandlekar2023; RoboGen2023; RoCoDA2024; Jiang2025DexMimicGen). Although these methods enrich demonstration sets, they remain constrained to variations of the same underlying tasks without expanding into new compositions.
@@
Another orthogonal direction emphasizes visual augmentation, including render-driven and vision-only pipelines that procedurally generate synthetic video datasets (Bonetto2023; Singh2024; Yu2024VisualParkour; Re3Sim2024; Real2Render2024), as well as generative and diffusion-based methods that augment images while holding actions constant (chen2023genaug; yu2023scaling). While effective for increasing visual diversity, these approaches do not provide transition-level data reflecting novel task semantics.
@@ Acknowledgments
DSB acknowledges support from the Center for Curiosity. MH and EE were partially supported by the Army Research Office under MURI award W911NF201- 0080 and DARPA award HR00112420305. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the view of DARPA, the US Army, or the US government.